OpenStudy (anonymous):

Word Problem: Solve by modeling the situation with a linear system. At a recent production of A Comedy of Errors the Community Theater brought in a total of $30,495 in revenue. If adult tickets were $9 and children's tickets were $6.50 how many tickets of each type were sold if 3800 tickets in all were sold?

OpenStudy (anonymous):

c=childs tickets a=adults tickets you're given the total sold so a+c=3800 and the total revenue plus price for each ticket 9a+6.50c=30495 now you can solve by substitution or elimination

OpenStudy (anonymous):

so im not sure how to set this up.. a + c = 3800 9a + 6.50c = 30495 I use elimination and I get something totally wrong. multiple the top by -9 for example, eliminate the a's.